OpenStudy (anonymous):

SOLVE FOR X: 1/2x = 20

OpenStudy (anonymous):

1=20 x 2x 1=40x 40/1=x 40??

OpenStudy (aroub):

x=40

OpenStudy (anonymous):

i might be completely wrong!!

OpenStudy (anonymous):

close >.< its x = 1/40

OpenStudy (anonymous):

oh right! thanks

OpenStudy (dumbcow):

\[\frac{1}{2x} = 20 \rightarrow 1 = 20(2x) \rightarrow 1 = 40x \rightarrow x = \frac{1}{40}\]

OpenStudy (anonymous):

i think they meant (1/2)x so it would b 40

OpenStudy (anonymous):

you guys had the right steps up to 1 = 40x. you divide both sides by 40 and get x = 1/40

OpenStudy (anonymous):

Is it (1/2) x or 1/(2x)?

OpenStudy (anonymous):

ah, good call, not really sure what the problem is.